# POLO - Point-based, multi-class animal detection

### v1

@@ 4.1.1 Dataset
We evaluate and test all models in this study on a publicly available set of drone images [17] from the Izembek lagoon in Alaska hosted on the LILA BC data repository. The dataset includes 9,267 images of waterfowl of size $8688\times 5792$ pixels, and a total of 521,270 pseudo-boxes for the classes “Brant goose” (424,790 boxes), “Canada goose” (47,561 boxes), “Gull” (5,631 boxes), “Emperor goose” (2,013 boxes), and “Other” (5,631 boxes). The data was divided into a training (80%), validation (5%), and test set (15%). Images were split into $640\times 640$ patches with 10% overlap, to match the input size expected by YOLOv8 and POLO. 95% of patches that did not contain animals were discarded. The remaining 5% were used as negative samples to reduce false positive detections. Bounding boxes that span multiple patches were clipped to the patch limits if at least 15% of the area of the box lied within the patch in question.

### v2

@@ 4.1.1 Dataset
We evaluate and test all models in this study on a publicly available set of drone images [17] from the Izembek lagoon in Alaska hosted on the LILA BC data repository. The dataset includes 9,2672 images of waterfowl of size $8688\times 5792$ pixels, and a total of 521,270 pseudo-boxes for the classes “Brant goose” (424,790 boxes), “Canada goose” (47,561 boxes), “Gull” (5,631 boxes), “Emperor goose” (2,013 boxes), and “Other” (5,631 boxes). The data was divided into a training (80%), validation (5%), and test set (15%). Images were split into $640\times 640$ patches with 10% overlap, to match the input size expected by YOLOv8 and POLO. 95% of patches that did not contain animals were discarded. The remaining 5% were used as negative samples to reduce false positive detections. Bounding boxes that span multiple patches were clipped to the patch limits if at least 15% of the area of the box lied within the patch in question.
@@ 6 Conclusion
Funding Information
This research has been carried out as part of the project WildDrone, funded by the European Union’s Horizon Europe Research and Innovation Program under the Marie Skłodowska-Curie Grant Agreement No. 101071224, the EPSRC funded Autonomous Drones for Nature Conservation Missions grant (EP/X029077/1), and the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number 22.00280.